A Complete Bibliography of Publications in the SIAM Journal on Applied Algebra and Geometry

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

26 August 2024 Version 1.08

Title word cross-reference

2 [LW22]. 3 [DKPR22, MSS22]. C^{2} [CGY17]. C^{3} [HL23]. j [Kos22]. C^{∗} [MMW21]. F_{q} [MSS22]. R^{3} [HL23]. Z^{n} [YMV21]. μ [SG17]. n [GO19a]. p [KL21]. q [Jan23]. X [CQU17].

-Action [MMW21]. -Adic [KL21]. -Bases [SG17]. -Fold [MSS22]. -Invariants [Kos22]. -Matroid [Jan23]. -Parameter [LW22]. -Rank [CQU17]. -Zeros [MSS22].

3-Periodic [HE24].

Abstract [Fri23]. Acting [HMR24]. Action [MMW21]. Actions [MKSV20]. Adapted [HHS21]. Adic [KL21]. AG [BACQ23].

Alexander [GK17]. Algebra [DLW21, MKSV20]. Algebraic [AP23, BKM21, CMS17, CLM24, DEW20, DEE^{+}23, DFR^{+}24, DM22, FO24, FdW22, KMZ21, KRV20, LS19, MS23, Sch18a, Sta21, TGND21]. Algorithm [Mic19, Sta21]. Algorithms [AKRS21, AKRS24b, CFW21, LMR23]. Alternatives [AM19]. Ambiguity [Edi19]. Analog [KW24]. Analogues [ACLN21]. Anomalies [SG17]. Application [CGY17, CDM^{+}24]. Applications [BDFHH20, Der18, DGHP23, KK17, MS19, NPSZ24]. Approach [BYF^{+}19, CMS21, CI22]. Approaches [AC22, Ave19]. Approximation [AP23, BDFHH20]. Arbitrary [ACLN21, BK18]. Arithmetic [Kop17].

Arrangements [GL22]. Aspects [Sta21]. Associated [SG17]. Atomically [YMV21]. Attacks [CLS17]. Auctions [JKS22]. Augmented [CKMW21]. Automorphisms [IL23]. Autonomy [SR18, Sha19]. Auxetic [BS17]. Averages [NW23]. Avoidance [GC24]. Axiomatic [CMS21].

Bad [SG17]. Barrier [ABGJ18]. Barycenter [CS22, CDM^{+}24]. Base [SG17]. Based [BYF^{+}19, Sta21, Tra21, Pol22]. Bases [KM19, SG17]. Bayesian [BSST23, Nic23]. Bent [GL22]. Betti [LW22, OS24]. Bigraded [LW22]. Bijectivity [MHR19]. Bilinear [KK17]. Biomembrane [CGY17]. Birationally [CDM18]. Bivariate [DY24]. Blum [Sta21]. Bodies [FE18]. Body [GO19a]. Border [CGJ19, LM17]. Bottleneck [DEW20]. Bound [DV21, FE18, HM22]. Boundary [FdW22, LM20]. Bounds [BS24, MRZ22]. Branching [BCHP18]. Breaking [BCGO17]. Buildings [Hir19].

Calibrated [Kil17]. Can [BK20, Shi22]. Carathéodory [LR23]. Cartan [ÖSS18]. Case [BMN22]. CAT [HH21]. Category [KW24]. Cauchy [El 20]. Cautionary [MW23]. Cells [BK18]. Central [BMN22, DH23]. Certifying [Sau19]. Characteristic [Mic19]. Characterization [AD23, NFB19]. Chief [Stu17]. Chiral [PST22]. Chordal [CP17]. Circuit [DIdW17, MS23, Wan22]. Circular [For23]. Classes [KE22]. Close [CLR24]. Clubs [NPSZ24]. Cluster [FO24]. Clustering [CMS21]. Code [CGJ^{+}17]. Coded [FHGHK17]. Codes [ACLN21, BH24, BACQ23, BNRS19, CLM24, CFS19, GR22, Jef20, Jef22, LS18b, MSS22]. Cohomology [YL17]. Colluding [FHGHK17]. Colored [CMMS23]. Combinatorial [Rav22]. Comeagre [HL23]. Common [GR22].

Communications [BDFHH20]. Comon [Dra24, Shi18]. Compact [HL23, HMR24]. Compatibility [BN20]. Complements [GR22]. Complete [HL23, KL23, MMW21]. Completely [KSZ24]. Completion [KKKR17]. Complex [HN23, Shi22]. Complexes [BHS^{+}19, MHP18]. Complexity [DIP20, KW24]. Computation [DFR^{+}24, LS19, YMV21]. Computer [LV20]. Computing [CS22, DEE^{+}23, HH21, Hir19, KMZ21, LM20, LW22, MS23, MW23]. Condition [Bür17]. Conditions [KK17]. Cone [FdW22, Sch21]. Cones [IL23]. Configurations [DH23]. Conformal [CS22]. Conjecture [Dra24, Shi18, Tra21]. Conley [MW23]. Connection [BYF^{+}19]. Connection-Based [BYF^{+}19]. Connections [MKSV20]. Connectivity [CMS17]. Consistency [MHP18]. Constants [BN20]. Constrained [MHP18]. Constraints [KMZ21]. Construction [BH24]. Containment [FNT17]. Context [ADV24]. Context-Specific [ADV24]. Contraction [ÖSS18]. Control [BCGO17]. Controls [MHP18]. Convergence [BMN22]. Convex [BKM21, CGJ^{+}17, El 20, FE18, HH21, Hir19, Sch18a, Sch21]. Convexity [CFS19]. Convolutional [KMMT22, KMST24]. Copositive [KSZ24]. Corrigendum [AKRS24b, Sha19]. Cost [BCGO17]. Counterexample [Dra24, Shi18]. Counting [CDM18, CHHM18]. Coupled [FO24]. Coupling [CHHM18]. Critical [KMST24]. Cryo [YL17]. Cryo-Electron [YL17]. Curvature [BKM21]. Curves [CCW19, KRV20, Kos22, Sch18a]. Cycle [Oba18]. Cycles [CCW19].

D [DKPR22]. Data [TGND21]. Databases [FHGHK17]. Datasets [NW23]. Decidability [CFS19]. Decomposable [ADV24]. Decomposing [DKPR22].

Decomposition [BHS^{+}19, CQU17, LMR23]. Decompositions [LM17]. Deconvolution [KK17]. Degree [BS24, BMN22, CLM24, DEW20, Hir19, LNRW23, MRW20, MMW21, RW17, SR18, Sha19]. del [BH24]. Denoising [Der18]. Density [Kra23]. Derivative [IL23]. Description [HMR24]. Design [BS17, DFR^{+}24]. Determinant [GO19a]. Determinantal [BFMT21, BvDD^{+}17, Vol19]. Determinants [Hir19]. Determined [Tur22]. Diagrams [GL23, MP22]. Differential [AGK^{+}24, Cra19, DGHP23, KRV20]. Dimension [LON24, MM17]. Dimensional [Edi19, FNT17]. Dimensions [KOT23]. Diophantine [ADOO17]. Dirac [KW20]. Directed [CMS21]. Directions [DJ23]. Discrete [AA19, HH21, Hir19]. Distance [ALR19, LON24, MRW20, MP22, NFB19]. Distinguishing [GL18]. Distributed [DLW21]. Distributions [AA19, OPW20]. Divisors [CLM24]. DLP [Mic19]. Do [BSV17]. Doubly [Dym18]. Dowling [Rav22]. Dream [BSV17]. Drone [BK20]. DSOS [AM19]. Dual [BACQ23]. Duality [GK17]. Dynamical [Cra19, DGHP23].

Edge [DY24]. Edit [MP22]. Editor [Stu17]. Editor-in-Chief [Stu17]. Eigenconfigurations [Koz18]. Eigenschemes [BGV21]. Eigenvalues [Bre17, KOT23, Kra23]. Eight [KOT23]. Elder [CKMW21]. Elder-Rule-Staircodes [CKMW21]. Electrical [For23]. Electron [YL17]. Elimination [DGHP23]. Elliptic [CCW19, Kos22]. Embeddings [HSW18]. Encodings [Kos22]. Enumerating [KE22]. Equation [RS21]. Equations [ADOO17, AGK^{+}24]. Equilibria [CDM18, CHHM18]. Equivalence [LS18a]. Equivariant [CNN20]. Erratum [Dra24, Jef22]. Errors [CLS17]. Estimation [AKRS21, AKRS24b, DM21, LON24, MHP18]. Estimator [AGK^{+}24]. Euclidean [HSW18, Hir19, MRW20, NFB19]. Evasion [GK17]. Evolution [LRR20]. Exact [Dym18]. Exceed [Shi22]. Exceptional [HSW18]. Existence [PST22]. Expected [Bre17]. experiment [OPPS22]. Experimental [DFR^{+}24]. Explicit [CLM24, HMR24, NFB19]. Exponential [MHR19]. Exponential/Generalized [MHR19]. Expressiveness [GL22]. Extension [BdSN17]. Extensions [ACLN21, CLM24, KMZ21]. Extraneous [SG17].

Facet [DJ23]. Factorization [CKM^{+}17]. Factors [SG17]. False [Tra21]. Families [BFMT21, MHR19]. Fan [MS19]. Fano [IZ17]. Feature [DEE^{+}23]. Fewnomials [BETC19]. Fibers [BBCY20]. Fields [Mic19]. Finite [CMS17, FNT17, Mic19, Tra21]. Finite-Dimensional [FNT17]. Finiteness [DH23]. Fixed [DH23, DJ23]. Flags [BACQ23]. Flat [KOT23]. Flexible [CGY17]. Fold [MSS22]. Forced [Ber22]. Forms [El 20, KL23]. Formulas [BFMT21]. Formulation [FO24]. Foundations [LRR20, MKSV20]. Frames [CMS17, CI22]. Frameworks [HE24]. Free [BSV17, CDM^{+}24, GL23]. Friendly [CCW19]. Fully [Koz18]. Function [KMST24]. Functions [AA19, BDFHH20, BCGO17, DEFT22, OPW20, OS24].

Galois [ACLN21, DS20, DKPR22]. Galois/Monodromy [DKPR22]. Gap [CGZ23]. Gaussian [AA19, AGK^{+}24, Bre17, CMMS23, DGS24]. General [CM20, Der18, JPRS23]. Generalized [El 20, JKS22, MHR19]. Generated [IL23, YMV21]. Generator [Oba18]. Generic [Ber22, SR18, Sha19]. Genetic [LRR20]. Genomic [ALR19]. Geometric [AC22, CLM24, DEE^{+}23, DIP20, HE24, NW23]. Geometries [Rav22]. Geometry [BKM21, BDLS23, DGS24, Edi19, FO24, KKKR17, KM19, KMMT22, Kop17, KL21, LS19, LM17]. Global [NFB19]. Good [BH24]. Gordan [LR23]. Gough [CM20, HSW18]. Graded [BBCY20]. Gram [HHS21]. Graph [CGG^{+}18, Pol22]. Graph-based [Pol22]. Graphical [CMMS23]. Graphs [AD23, Kra23]. Group [HSW18]. Groups [CDM^{+}24, DKPR22, HMR24, LS19, ÖSS18]. Guided [KE22].

Harmonic [BC24]. Hausdorff [LM20]. Hear [BK20]. Hierarchical [CMS21]. Higher [BdSN17, HKU17, KM19, Rav22]. Higher-Order [HKU17]. Higher-Weight [Rav22]. Hilbert [HN23, LV20, OS24]. Hilbert-Samuel [HN23]. Homogeneous [Las19]. Homology [AC22, BK23, BC24, HOST19, LW22, MKPGC19, Oba18]. Hulls [CLM24, Sch18a]. Hyperbolic [Sau19]. Hyperbolicity [IL23]. Hypercube [BS24]. Hyperplane [GL22].

Ideals [CP17, CMMS23, KW20]. Identifiability [CQU17, DGHP23, KK17, OPPS22]. Improvements [Sta21]. Including [Sta21]. Inclusion [BN20]. Inclusions [Cra19]. Independence [RW17]. Index [CDM18, MW23]. Inequalities [Ave19, El 20]. Inference [ALR19]. Infinitesimal [Ber22]. Infinity [SG17]. Information [DS20, FHGHK17]. Injectivity [KK17]. Innovation [LRR20]. Integer [OPW20]. Integral [KL21]. Integrals [BSST23, CDM^{+}24]. Integration [MKSV20]. Interior [ABGJ18, DY24]. Interpolation [BdSN17]. Intersecting [Bür17]. Intersection [CDM18]. Invariant [AKRS21, CDM18, CI22, MKSV20, AKRS24b]. Invariants [HL23, Kos22, LS19]. Inverse [KK17]. Irrational [CBG^{+}23].

Irrationality [CKM^{+}17]. Irreducibility [CMS17]. Isometries [YMV21]. Isometry [BACQ23]. Isometry-Dual [BACQ23]. Isoperimetric [DLSS19]. Isotopy [KE22]. Isotropy [CI22]. Iterated [CDM^{+}24, DEFT22]. Iterated-Integrals [CDM^{+}24]. Iterated-Sums [DEFT22].

Jewel [BN20]. Joint [PSW21].

Kalai [Sta21]. Kernel [AP23]. Khovanskii [KM19]. Kippenhahn [PSW21]. Koszul [BFMT21]. Koszul-Type [BFMT21]. Kronecker [MM17]. Kummer [CLM24]. Kuramoto [CDM18, CHHM18].

Laman [CGG^{+}18]. Laplacian [KOT23]. Large [KOT23]. Lattice [LS18a]. Lattices [BK18, MP22, Rav22]. LCD [CLM24]. Leaf [GO19b]. Learning [BYF^{+}19, DJ23]. Lemmas [MS19]. Lie [CDM^{+}24, MKSV20]. Likelihood [AKRS21, AKRS24b, AGK^{+}24, DM21, DGS24, LNRW23, MMW21, RW17]. Line [BRST23]. Linear [ADOO17, Ave19, Bür17, GR22, KMZ21, KMMT22, KMST24, RS21]. Links [HN23]. Lipschitz [GL23]. Lipschitz-Free [GL23]. Local [CFS19, KW20]. Locating [CHHM18]. Locus [CM20]. Log [ABGJ18, ALR19]. Log-Barrier [ABGJ18]. Log-Det [ALR19]. Lower [DV21, FE18]. LWE [Sta21].

Makes [CGJ^{+}17]. Manifold [BYF^{+}19]. Manifolds [DEE^{+}23, GC24, MKSV20]. Many [BSS18, BACQ23]. Many-Point [BACQ23]. Maps [BBCY20, CQU17, DS20, MHR19]. Marginals [MM17]. Mathematical [LRR20]. Matrices [NFB19]. Matrix [Ave19, BN20, CI22, CKM^{+}17, DM21, GG21, GO19a, KSZ24, LM17]. Matroid [Jan23, MS23, Tra21]. Matroid-Based [Tra21]. Maximum [AKRS21, AKRS24b, AGK^{+}24, DM21, LNRW23, MMW21, RW17]. Maxout [MRZ22]. Measure [LM20]. Measurements [BN20]. Message [Stu17]. Method [BSS18, LS19]. Methods [ABGJ18, TGND21]. Metric [AAF18, ACLN21, BNRS19, CKMW21]. Metrics [CR20]. Microscopy [YL17]. Minimal [DKPR22, Kil17, KE22, LW22]. Minimization [NFB19]. Minkowski [LR23, MRZ22]. Mixed [BFMT21]. Mixtures [KW20, RW17]. Model [CDM18, CHHM18, JPRS23, OPPS22]. Modeling [HE24]. Modelling [CGY17]. Models [ADV24, ALR24, AGK^{+}24, CMMS23, DM21, DGHP23, MM17, RS21, RW17]. Modules [BdSN17]. Moment [DS20, KW20]. Monodromy [DKPR22]. Monomials [HM22]. Monotone [BDLS23]. Monotonicity [KL23]. Morphisms [Jef20, Jef22]. Morse [MHP18]. Motion [CFW21, ÖSS18]. MRD [GR22]. Muller [ACLN21]. Multi [BBCY20, OPPS22, RS21]. Multi-experiment [OPPS22]. Multi-Graded [BBCY20]. Multi-Trek [RS21]. Multigraded [OS24]. Multilabeled [MS19]. Multilinear [BFMT21, DLW21]. Multiparameter [HOST19]. Multiplication [LM17]. Multiplicative [CGJ19]. Multiplicities [HN23]. Multiplicity [DIP20]. Multitriangulations [RS24]. Multiview [BRST23, MRW20].

nc [HH21]. nc-Rank [HH21]. Networks [CMS21, CP17, Cra19, DP17, For23, GL22, GL18, KMMT22, KMST24, MRZ22, Nic23]. Neural [CFS19, CGJ^{+}17, GL22, Jef20, Jef22]. Nilpotent [CDM^{+}24]. Non [CLM24, Nic23]. Non-special [CLM24]. Non-toric [Nic23]. Noncommutative [Vol19]. Nonnegative [CKM^{+}17, DIdW17, KL23, Wan22]. Nonnegativity [Sau19]. Norm [CMS17]. Normal [DM21]. Norms [BK18, Nie17]. Nuclear [Nie17]. Number [Bre17, BETC19, CGG^{+}18, LS18b, MRZ22]. Numbers [LW22, OS24, Rav22]. Numerical [LS19, MKSV20, PSW21].

Obstacle [GC24]. Obstruction [CFS19]. Obstructions [DIP20]. Occurrence [DIP20]. ODEs [OPPS22]. One [CLR24, CR20, CHHM18, Edi19, IL23, KKKR17]. One-Dimensional [Edi19]. onto [BBCY20]. Operator [FNT17]. Operators [BKM21]. Optimal [AAF18, Ave19, BCGO17, JKS22, KK17, Kos22, Oba18]. Optimization [AM19, Ave19, BMN22, DM22, HH21, HE24, Hir19, OPW20, Sau19]. Optimize [Pol22]. Orbit [HMR24, LS18a, YMV21]. Orbits [DS20]. Order [HKU17, Sch21]. Ordinary [GLSV19]. Orientable [HL23]. Orthogonal [BBCY20]. Other [LM17].

Packings [FJM18]. Pairing [CCW19]. Pairing-Friendly [CCW19]. Parallel [BYF^{+}19]. Parameter [CR20, LW22, OPPS22]. Parametric [OPW20]. Parametrized [CFW21]. Partial [GO19b]. Partitioned [NW23]. Path [BMN22]. Paths [BDLS23, GL23]. Paulsen [CI22]. Perfect [AD23]. Periodic [BS17, HE24, KE22, Kra23]. Permanent [Shi21]. Permutahedra [JKS22]. Persistence [BdSN17, CR20, GL23, MP22]. Persistent [AC22, BK23, BC24, HOST19, LW22, MKPGC19, Oba18]. Perspectives [DM22]. Perturbation [HKU17]. Pezzo [BH24]. Pfaffians [RS24]. Phase [Edi19]. Phaseless [GG21]. Phylogenetic [DP17, For23, GO19b, GL18]. Phylogenomic [ALR24]. Picard [LS19, LS18b]. Pivot [BDLS23]. Planar [For23]. Plane [Sch21]. Planning [CFW21]. Platform [CM20]. Platforms [HSW18].

Point [ABGJ18, BACQ23]. Points [KMST24, SG17]. Polyhedral [BSS18, BDLS23]. Polynomial [ABGJ18, Ave19, CP17, CQU17, Cra19, HMR24, KMZ21, LNRW23, Mic19, MHR19, Pol22, Sau19]. Polynomials [DIdW17, LMR23, Las19, MS23, MSS22, Mic19, Vol19, Wan22]. Polytope [AKRS24a]. Polytopes [BCHP18, DJ23, LS18a]. Posets [KMZ21]. Positive [BSS18, FE18, GK17, KSZ24]. Positivstellensatz [BS24, DIdW17]. Preordered [Fri23]. Presentations [LW22]. Principles [BS17]. Private [FHGHK17]. Probability [CLR24, KSZ24]. Problem [CLS17, CI22]. Problems [BCGO17, DKPR22, GC24, KK17, Kil17]. Product [CGJ19, MM17, SG17]. Profiles [DLSS19]. Programming [BS21]. Projection [BBCY20]. Projections [DGHP23]. Projective [Bür17, DGS24, FJM18]. Projectivization [Jan23]. Propagation [JPRS23]. Pursuit [GK17]. Putinar [BS24].

Quadrics [MMW21]. Quantifying [LRR20]. Quantum [BN20, CBG^{+}23, DS20, PSW21]. Quasi [Mic19]. Quasi-Polynomial [Mic19]. Quasisymmetric [DEFT22]. Queueing [Pol22]. Quiver [CI22, DM21].

Radial [CI22]. Random [BETC19, CLR24, KSZ24]. Ranges [PSW21]. Rank [ACLN21, BNRS19, CLR24, CGJ19, CQU17, CHHM18, FE18, GG21, HH21, HM22, IL23, KKKR17, KM19, LM17, Shi21]. Rank-Metric [ACLN21, BNRS19]. Rank-One [CLR24, CHHM18, IL23, KKKR17]. Rate [AKRS24a, BMN22]. Rational [AGK^{+}24, BBCY20, SG17]. Rationality [CM20]. Reaction [Cra19]. Real [Bre17, BETC19, FJM18, Koz18, NFB19, Sch18a]. Realization [FNT17]. Realizations [CGG^{+}18]. Reconstruction [AAF18, CNN20, DKPR22, GLSV19]. Reconstructions [PST22]. Recovery [Dym18]. Reduction [GC24, Pol22]. Reed [ACLN21]. Region [AKRS24a]. Regions [MRZ22]. Related [LS19, OPW20]. Relaxation [Dym18]. Relaxations [IL23]. ReLU [GL22]. Representation [BNRS19, El 20, Sch18a, Sch21]. Representations [BvDD^{+}17, DM21, GO19b, Vol19]. Representative [Oba18]. Requires [CKM^{+}17]. Resolutions [BSV17]. Resonant [Kop17]. Respect [BK18]. Reticulate [LRR20]. Retrieval [Edi19, FHGHK17]. Riemannian [GC24, HE24]. Rigidity [Ber22, MS23]. Ring [Sta21, Sta21]. Ring-Based [Sta21]. Ring-LWE [Sta21]. RLWE [CLS17]. RNA [BCHP18]. Robust [CMS21, HE24]. Room [BK20]. Rossby [Kop17]. Rotations [Ber22]. Rule [CKMW21]. Rules [BDLS23].

Samuel [HN23]. Scaling [AKRS21, AKRS24b]. Scheme [CGY17]. Schemes [IZ17, LV20]. Schwarz [El 20]. SDSOS [AM19]. SE [HL23]. Search [CLS17]. Second [Sch21]. Second-Order [Sch21]. Segrè [HSW18]. Selection [Mic19]. Semialgebraic [BK23, LM20]. Semidefinite [AM19, Ave19, BMN22, BS21, FE18, Sch18a]. Semifields [Fri23]. Semirings [Fri23]. Separation [RS21]. Sequences [ALR19]. Series [DEFT22]. Servers [FHGHK17]. Service [AKRS24a]. Set [HL23]. Sets [BK23, GO19b, KL23, Las19, LM20, Sch21]. Shadows [Sch18b]. Shape [BK20]. Sharp [MRZ22]. Signal [Der18]. Signatures [CDM^{+}24, DEFT22, GL23, KRV20]. Signomial [DM22]. Silhouettes [GLSV19]. Simplicial [CNN20]. Single [DY24].

Singular [BHS^{+}19, Der18, HKU17, Tur22]. Singularities [CM20, GLSV19]. Size [Ave19]. Sizes [DEE^{+}23]. Slope [JPRS23]. Smale [MHP18]. Small [CLM24, CLS17, LS18b, Mic19]. Solutions [ADOO17, BSS18]. Solving [Sta21]. Sonc [FdW22]. Sonc-Cone [FdW22]. Space [DP17, GO19b, KMST24, LON24]. Spaces [CKMW21, FJM18, GL23, HH21, HMR24]. Sparse [ADOO17, BSS18, BS21, LNRW23, MSS22]. Sparseness [GR22]. Spatial [DH23]. Special [HSW18, CLM24]. Species [ALR19]. Specific [ADV24]. Spectrahedra [HHS21]. Spectrahedral [FNT17, Sch18b]. Sperner [MS19]. Sphere [CGY17]. Splines [DV21, DY24]. Split [For23]. Square [NFB19]. Squares [AM19, AD23, BSV17, BS21, El 20]. Stability [OS24]. Stabilization [CR20]. Stable [Vol19]. Staircodes [CKMW21]. Stars [DV21]. States [PSW21]. Statistical [AGK^{+}24, MHP18]. Statistics [MKPGC19]. Stewart [CM20, HSW18]. Stochastic [Dym18]. Storage [DLW21]. Stratifying [HOST19]. Strong [SG17]. Stronger [DIP20]. Strongly [ABGJ18]. Structural [DGHP23, RS21]. Structure [BCHP18]. Study [LRR20]. Subconfigurations [DH23]. Subdivision [CGY17]. Subgroups [YMV21]. Subject [KMZ21]. Sublevel [KL23, Las19]. Subrank [CGZ23, Shi22]. Subspace [Bür17]. Subspaces [GR22]. Sufficient [MKPGC19]. Sum [AM19, AD23, El 20]. Sums [BSV17, BS21, DEFT22, DIdW17, MRZ22]. Surfaces [BH24, BBCY20, GLSV19, HL23, KE22, LS18b, SG17]. Symmetric [CLR24, NFB19, Nie17, Shi22]. Symmetrically [CMMS23]. Symmetries [ALR24, Dym18]. Symmetrized [NFB19]. Symmetry [Ber22, BCGO17, GC24, HHS21, KE22, LM17, LR23, Pol22]. Symmetry-Forced [Ber22].

Synchronization [ÖSS18]. Systems [BFMT21, BSS18, Cra19, For23, FNT17, LNRW23].

Tale [MW23]. Tangent [LON24].

Template [MHP18]. Tensegrity [HE24]. Tensor [Bre17, BNRS19, CLR24, CGJ19, KKKR17, Nie17, SG17, Shi22, TGND21]. Tensors [BGV21, CGZ23, Koz18, LMR23, LM17, Tur22]. Ternary [BGV21]. Tetrahedral [DV21]. Their [GLSV19, IL23, NPSZ24, Tur22, Vol19]. Theorem [PSW21]. Theorems [LR23]. Theoretic [CI22]. Theory [AKRS21, AKRS24b, Der18, DFR^{+}24, DS20, DIP20, FO24, OPPS22, Pol22]. Theta [AA19, BDFHH20]. Tight [CMS17]. Tightest [Oba18]. Tilings [KE22]. Time [DEFT22, Mic19]. Topological [GL22, LRR20]. Topologically [MHP18]. Topology [CFW21]. Tori [HMR24, KOT23]. Toric [BSST23, CMMS23, Cra19, IZ17, MSS22, Nic23]. Total [DLSS19]. Totally [DY24]. Tractable [AM19]. Transcendental [LS19]. Transformations [FN18]. Translations [Ber22]. Transport [AAF18, BYF^{+}19]. Transversality [GL22]. Tree [ALR19, ALR24, GO19b]. Trek [RS21]. Triads [FN18, Kop17]. Triangulation [DY24]. Trifocal [Kil17]. Triply [KE22]. Trivariate [MSS22]. Tropical [DEFT22, KM19, MKPGC19, RS24]. Tuples [Tur22]. Two [LV20, PST22]. Type [BFMT21].

Unfolding [BYF^{+}19]. Unified [HMR24]. Uniform [BvDD^{+}17]. Unit [CMS17]. Using [ALR19, CDM18, HE24].

Valuation [Tra21]. Valuations [KM19]. Value [BHS^{+}19]. Values [Der18, HKU17]. Vanishing [CMMS23]. Variation [DLSS19]. Varieties [AP23, BSST23, BRST23, CMS17, DEW20, DGS24, IZ17]. Variety [Bür17, Kil17, MRW20]. Varying [Bür17]. Vergleichsstellensätze [Fri23]. Versions [MS19]. Vertex [DV21]. Vertices [MRZ22]. Via [AAF18, AA19, DM21, DGHP23, HH21, Hir19, MHP18, ÖSS18, Sau19]. View [PST22]. Vision [LV20]. Voicing [FN18]. Volume [Las19, Oba18]. Volume-Optimal [Oba18]. Volumes [KL23]. Voronoi [BK18].

Walks [CBG^{+}23]. Waring [HM22, Shi21]. Wasserman [Sta21]. Wasserstein [LON24]. Wave [Kop17]. Weak [BH24]. Weight [Rav22]. Weyl [HMR24, LR23]. Whitney [Rav22]. Wildfire [JPRS23]. Wind [JPRS23]. Worst [BMN22]. Worst-Case [BMN22].

x [Shi21].

Zeros [BETC19, MSS22].

References

Agostini:2019:DGD

[AA19] Daniele Agostini and Carlos Améndola. Discrete Gaussian distributions via theta functions. SIAM Journal on Applied Algebra and Geometry, 3(1):1–30, ???? 2019. CODEN SJAABQ. ISSN 2470-6566.

Adamaszek:2018:MRO

[AAF18] Michał Adamaszek, Henry Adams, and Florian Frick. Metric reconstruction via optimal transport. SIAM Journal on Applied Algebra and Geometry, 2 (4):597–619, ???? 2018. CO-DEN SJAABQ. ISSN 2470-

Allamigeon:2018:LBI

[ABGJ18] Xavier Allamigeon, Pascal Benchimol, Stéphane Gaubert, and Michael Joswig. Log-barrier interior point methods are not strongly polynomial. SIAM Journal on Applied Algebra and Geometry, 2(1):140–178, ???? 2018. CODEN SJAABQ. ISSN 2470-6566.

Adams:2022:GAP

[AC22] Henry Adams and Baris Coskunuzer. Geometric approaches to persistent homology. SIAM Journal on Applied Algebra and Geometry, 6(4):685–710, ???? 2022. CODEN SJAABQ. ISSN 2470-

6566. URL https://epubs.
siam.org/doi/10.1137/21M1422914.

Augot:2021:RMC

[ACLN21] Daniel Augot, Alain Couvreur, Julien Lavauzelle, and Alessandro Neri. Rankmetric codes over arbitrary Galois extensions and rank analogues of Reed–Muller codes. SIAM Journal on Applied Algebra and Geometry, 5(2):165– 199, ???? 2021. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/20M1348583.

Ahmadi:2023:SSC

[AD23] Amir Ali Ahmadi and Cemil Dibek. A sum of squares characterization of perfect graphs. SIAM Journal on Applied Algebra and Geometry, 7(4):685– 715, ???? 2023. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/22M1530410.

Aliev:2017:SSL

[ADOO17] Iskander Aliev, Jesús A. De Loera, Timm Oertel, and Christopher O’Neill. Sparse solutions of linear Diophantine equations. SIAM Journal on Applied Algebra and Geometry, 1(1):239–253, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Alexandr:2024:DCS

[ADV24] Yulia Alexandr, Eliana Duarte, and Julian Vill. Decomposable context-specific models. SIAM Journal on Applied Algebra and Geometry, 8(2):363–393, May 2024. ISSN 2470-6566.

Amendola:2024:DEG

[AGK^{+}24] Carlos Améndola, Lukas Gustafsson, Kathlén Kohn, Orlando Marigliano, and Anna Seigal. Differential equations for Gaussian statistical models with rational maximum likelihood estimator. SIAM Journal on Applied Algebra and Geometry, 8 (3):465–492, ???? 2024. CO-DEN SJAABQ. ISSN 2470-

6566. URL https://epubs.
siam.org/doi/10.1137/23M1569228.

Amendola:2021:ITS

[AKRS21] Carlos Améndola, Kathlén Kohn, Philipp Reichenbach, and Anna Seigal. Invariant theory and scaling algorithms for maximum likelihood estimation. SIAM Journal on Applied Algebra and Geometry, 5(2):304– 337, ???? 2021. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/20M1328932.

See corrigendum [AKRS24b].

Alfarano:2024:SRR

[AKRS24a] Gianira N. Alfarano, Altan B. Kiliç, Alberto Ravagnani, and Emina Soljanin. The service rate region polytope. SIAM Journal on Applied Algebra and Geometry, 8(3):553– 582, ???? 2024. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/23M1557829.

Amendola:2024:CIT

[AKRS24b] Carlos Améndola, Kathlén Kohn, Philipp Reichenbach, and Anna Seigal. Corrigendum: Invariant theory and scaling algorithms for maximum likelihood estimation. SIAM Journal on Applied Algebra and Geometry, 8(2):460– 463, ???? 2024. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/23M1588317.

See [AKRS21].

Allman:2019:STI

[ALR19] Elizabeth S. Allman, Colby Long, and John A. Rhodes. Species tree inference from genomic sequences using the log-Det distance. SIAM Journal on Applied Algebra and Geometry, 3(1):107–127, ???? 2019. CODEN SJAABQ. ISSN 2470-6566.

Allman:2024:PMT

[ALR24] Elizabeth S. Allman, Colby Long, and John A. Rhodes. Phylogenomic models from tree symmetries. SIAM Journal on Applied Algebra and Geometry, 8(1):114–137, February 2024. ISSN 2470-6566.

Ahmadi:2019:DSO

[AM19] Amir Ali Ahmadi and Anirudha Majumdar. DSOS and SDSOS optimization: More tractable alternatives to sum of squares and semidefinite optimization. SIAM Journal on Applied Algebra and Geometry, 3(2):193–230, ???? 2019. CODEN SJAABQ. ISSN 2470-6566.

Altschuler:2023:KAA

[AP23] Jason M. Altschuler and Pablo A.

Parrilo. Kernel approximation on algebraic varieties. SIAM Journal on Applied Algebra and Geometry, 7(1):1– 28, ???? 2023. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/21M1425050.

Averkov:2019:OSL

[Ave19] Gennadiy Averkov. Optimal size of linear matrix inequalities in semidefinite approaches to polynomial optimization. SIAM Journal on Applied Algebra and Geometry, 3(1):128–151, ???? 2019. CODEN SJAABQ. ISSN 2470-6566.

Bras-Amoros:2023:IDF

[BACQ23] Maria Bras-Amorós, Alonso S. Castellanos, and Luciane Quoos. Isometry-dual flags of manypoint AG codes. SIAM Journal on Applied Algebra and Geometry, 7(4):786–808, November 2023. ISSN 2470-6566.

Botbol:2020:FMG

[BBCY20] Nicolás Botbol, Laurent Busé, Marc Chardin, and Fatmanur Yildirim. Fibers of multi-graded rational maps and orthogonal projection onto rational surfaces. SIAM Journal on Applied Algebra and Geometry, 4 (2):322–353, ???? 2020. CO-DEN SJAABQ. ISSN 2470-6566.

Basu:2024:HPH

[BC24] Saugata Basu and Nathanael Cox. Harmonic persistent homology. SIAM Journal on Applied Algebra and Geometry, 8 (1):189–224, March 2024. ISSN 2470-6566.

Bloch:2017:OCP

[BCGO17] Anthony M. Bloch, Leonardo J. Colombo, Rohit Gupta, and Tomoki Ohsawa. Optimal control problems with symmetry breaking cost functions. SIAM Journal on Applied Algebra and Geometry, 1(1):626–646, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Barrera-Cruz:2018:SRB

[BCHP18] Fidel Barrera-Cruz, Christine Heitsch, and Svetlana Poznanović. On the structure of RNA branching polytopes. SIAM Journal on Applied Algebra and Geometry, 2(3):444–461, ???? 2018. CODEN SJAABQ. ISSN 2470-6566.

Barreal:2020:ATF

[BDFHH20] Amaro Barreal, Mohamed Taoufiq Damir, Ragnar Freij-Hollanti, and Camilla Hollanti. An approximation of theta functions with applications to communications. SIAM Journal on Applied Algebra and Geometry, 4 (4):471–501, ???? 2020. CO-DEN SJAABQ. ISSN 2470-6566.

Black:2023:PGP

[BDLS23] Alexander E. Black, Jesús A.

De Loera, Niklas Lütjeharms, and Raman Sanyal. The polyhedral geometry of pivot rules and monotone paths. SIAM Journal on Applied Algebra and Geometry, 7(3):623– 650, ???? 2023. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/22M1475910.

Bubenik:2017:HIE

[BdSN17] Peter Bubenik, Vin de Silva, and Vidit Nanda. Higher interpolation and extension for persistence modules. SIAM Journal on Applied Algebra and Geometry, 1(1):272–284, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Bernstein:2022:GSF

[Ber22] Daniel Irving Bernstein. Generic symmetry-forced infinitesimal rigidity: Translations and rotations. SIAM Journal on Applied Algebra and Geometry, 6(2): 190–215, ???? 2022. CODEN SJAABQ. ISSN 2470-6566.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/20M1346961.

Burgisser:2019:NRZ

[BETC19] Peter Bürgisser, Alperen A. Ergür, and Josué Tonelli-Cueto. On the number of real zeros of random fewnomials. SIAM Journal on Applied Algebra and Geometry, 3(4):721–732, ???? 2019. CODEN SJAABQ. ISSN 2470-6566.

Bender:2021:KTD

[BFMT21] Matías R. Bender, Jean-Charles Faugère, Angelos Mantzaflaris, and Elias Tsigaridas. Koszultype determinantal formulas for families of mixed multilinear systems. SIAM Journal on Applied Algebra and Geometry, 5(4): 589–619, ???? 2021. CODEN SJAABQ. ISSN 2470-6566.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/20M1332190.

Beorchia:2021:ETT

[BGV21] Valentina Beorchia, Francesco Galuppi, and Lorenzo Venturello. Eigenschemes of ternary tensors. SIAM Journal on Applied Algebra and Geometry, 5 (4):620–650, ???? 2021. CO-DEN SJAABQ. ISSN 2470-

                                        6566. URL https://epubs.
                                        siam.org/doi/10.1137/20M1355410.

Blache:2024:CGC

[BH24] Régis Blache and Emmanuel Hallouin. Construction of good codes from weak del Pezzo surfaces. SIAM Journal on Applied Algebra and Geometry, 8(2): 414–459, ???? 2024. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/23M1549882.

Brake:2019:SVD

[BHS^{+}19] Danielle A. Brake, Jonathan D.

Hauenstein, Frank-Olaf Schreyer, Andrew J. Sommese, and Michael E. Stillman. Singular value decomposition of complexes. SIAM Journal on Applied Algebra and Geometry, 3 (3):507–522, ???? 2019. CO-DEN SJAABQ. ISSN 2470-6566.

Blomer:2018:VCL

[BK18] Johannes Blömer and Kathlén Kohn. Voronoi cells of lattices with respect to arbitrary norms. SIAM Journal on Applied Algebra and Geometry, 2(2):314–338, ???? 2018. CODEN SJAABQ. ISSN 2470-6566.

Boutin:2020:DCH

[BK20] Mireille Boutin and Gregor Kemper. A drone can hear the shape of a room. SIAM Journal on Applied Algebra and Geometry, 4(1):123–140, ???? 2020. CODEN SJAABQ. ISSN 2470-

Basu:2023:PHS

[BK23] Saugata Basu and Negin Karisani. Persistent homology of semialgebraic sets. SIAM Journal on Applied Algebra and Geometry, 7(3):651–684, ???? 2023. CO-DEN SJAABQ. ISSN 2470-

6566. URL https://epubs.
siam.org/doi/10.1137/22M1494415.

Bettiol:2021:CAG

[BKM21] Renato G. Bettiol, Mario Kummer, and Ricardo A. E. Mendes. Convex algebraic geometry of curvature operators. SIAM Journal on Applied Algebra and Geometry, 5(2):200– 228, ???? 2021. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/20M1350777.

Basu:2022:CPS

[BMN22] Saugata Basu and Ali Mohammad-Nezhad. On the central path of semidefinite optimization: Degree and worst-case convergence rate. SIAM Journal on Applied Algebra and Geometry, 6(2): 299–318, ???? 2022. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/21M1419933.

Bluhm:2020:CQM

[BN20] Andreas Bluhm and Ion Nechita. Compatibility of quantum measurements and inclusion constants for the matrix jewel. SIAM Journal on Applied Algebra and Geometry, 4(2):255–296, ???? 2020. CODEN SJAABQ. ISSN 2470-6566.

Byrne:2019:TRR

[BNRS19] Eimear Byrne, Alessandro Neri, Alberto Ravagnani, and John Sheekey. Tensor representation of rank-metric codes. SIAM Journal on Applied Algebra and Geometry, 3(4):614–643, ???? 2019. CODEN SJAABQ. ISSN 2470-6566.

Breiding:2017:ENE

[Bre17] Paul Breiding. The expected number of eigenvalues of a real Gaussian tensor. SIAM Journal on Applied Algebra and Geometry, 1(1):254–271, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Breiding:2023:LMV

[BRST23] Paul Breiding, Felix Rydell, Elima Shehu, and Angélica Torres. Line multiview varieties. SIAM Journal on Applied Algebra and Geometry, 7(2):470– 504, ???? 2023. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/22M1482263.

Borcea:2017:NPA

[BS17] Ciprian S. Borcea and Ileana Streinu. New principles for auxetic periodic design. SIAM Journal on Applied Algebra and Geometry, 1(1):442–458, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Blekherman:2021:SSS

[BS21] Grigoriy Blekherman and Kevin Shu. Sums of squares and sparse semidefinite programming. SIAM Journal on Applied Algebra and Geometry, 5(4): 651–674, ???? 2021. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/20M1376170.

Baldi:2024:DBP

[BS24] Lorenzo Baldi and Lucas Slot. Degree bounds for Putinar’s Positivstellensatz on the hypercube. SIAM Journal on Applied Algebra and Geometry, 8(1):1– 25, January 2024. ISSN 2470-6566.

Bihan:2018:PMS

[BSS18] Frédéric Bihan, Francisco Santos, and Pierre-Jean Spaenlehauer. A polyhedral method for sparse systems with many positive solutions. SIAM Journal on Applied Algebra and Geometry, 2(4):620–645, ???? 2018. CODEN SJAABQ. ISSN 2470-6566.

Borinsky:2023:BIT

[BSST23] Michael Borinsky, Anna-Laura Sattelberger, Bernd Sturmfels, and Simon Telen. Bayesian integrals on toric varieties. SIAM Journal on Applied Algebra and Geometry, 7(1):77– 103, ???? 2023. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/22M1490569.

Blekherman:2017:DSS

[BSV17] Grigoriy Blekherman, Rainer Sinn, and Mauricio Velasco. Do sums of squares dream of free resolutions? SIAM Journal on Applied Algebra and Geometry, 1(1):175–199, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Burgisser:2017:CIP

[Bür17] Peter Bürgisser. Condition of intersecting a projective variety with a varying linear subspace. SIAM Journal on Applied Algebra and Geometry, 1(1):111–125, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Boralevi:2017:UDR

[BvDD^{+}17] Ada Boralevi, Jasper van Doornmalen, Jan Draisma, Michiel E. Hochstenbach, and Bor Plestenjak. Uniform determinantal representations. SIAM Journal on Applied Algebra and Geometry, 1(1):415–441, ???? 2017. CO-DEN SJAABQ. ISSN 2470-6566.

Budninskiy:2019:PTU

[BYF^{+}19] Max Budninskiy, Gloria Yin, Leman Feng, Yiying Tong, and Mathieu Desbrun. Parallel transport unfolding: a connection-based manifold learning approach. SIAM Journal on Applied Algebra and Geometry, 3(2):266–291, ???? 2019. CODEN SJAABQ. ISSN 2470-6566.

Coutinho:2023:IQW

[CBG^{+}23] Gabriel Coutinho, Pedro Ferreira Baptista, Chris Godsil, Thomás Jung Spier, and Reinhard Werner. Irrational quantum walks. SIAM Journal on Applied Algebra and Geometry, 7(3):567–584, ???? 2023. CO-DEN SJAABQ. ISSN 2470-

6566. URL https://epubs.
siam.org/doi/10.1137/22M1521262.

Chiesa:2019:CPF

[CCW19] Alessandro Chiesa, Lynn Chua, and Matthew Weidner. On cycles of pairing-friendly elliptic curves. SIAM Journal on Applied Algebra and Geometry, 3 (2):175–192, ???? 2019. CO-DEN SJAABQ. ISSN 2470-6566.

Chen:2018:CEK

[CDM18] Tianran Chen, Robert Davis, and Dhagash Mehta. Counting equilibria of the Kuramoto model using birationally invariant intersection index. SIAM Journal on Applied Algebra and Geometry, 2(4):489–507, ???? 2018. CODEN SJAABQ. ISSN 2470-6566.

Clausel:2024:BFN

[CDM^{+}24] Marianne Clausel, Joscha Diehl, Raphael Mignot, Leonard Schmitz, Nozomi Sugiura, and Konstantin Usevich. The barycenter in free nilpotent Lie groups and its application to iteratedintegrals signatures. SIAM Journal on Applied Algebra and Geometry, 8(3):519–552, ???? 2024. CODEN SJAABQ. ISSN 2470-6566. URL https://

epubs.siam.org/doi/10.1137/
23M159024X.

Chen:2019:NCD

[CFS19] Aaron Chen, Florian Frick, and Anne Shiu. Neural codes, decidability, and a new local obstruction to convexity. SIAM Journal on Applied Algebra and Geometry, 3(1):44–66, ???? 2019. CODEN SJAABQ. ISSN 2470-6566.

Cohen:2021:TPM

[CFW21] Daniel C. Cohen, Michael Farber, and Shmuel Weinberger. Topology of parametrized motion planning algorithms. SIAM Journal on Applied Algebra and Geometry, 5(2):229–249, ???? 2021. CODEN SJAABQ. ISSN 2470-6566. URL https://

epubs.siam.org/doi/10.1137/
20M1358505.

Capco:2018:NRL

[CGG^{+}18] Jose Capco, Matteo Gallet, Georg Grasegger, Christoph Koutschan, Niels Lubbes, and Josef Schicho. The number of realizations of a Laman graph. SIAM Journal on Applied Algebra and Geometry, 2(1):94–125, ???? 2018. CODEN SJAABQ. ISSN 2470-6566.

Curto:2017:WMN

[CGJ^{+}17] Carina Curto, Elizabeth Gross, Jack Jeffries, Katherine Morrison, Mohamed Omar, Zvi Rosen, Anne Shiu, and Nora Youngs. What makes a neural code convex? SIAM Journal on Applied Algebra and Geometry, 1(1):222–238, ???? 2017.

CODEN SJAABQ. ISSN 2470-6566.

Christandl:2019:BRM

[CGJ19] Matthias Christandl, Fulvio Gesmundo, and Asger Kjærulff Jensen. Border rank is not multiplicative under the tensor product. SIAM Journal on Applied Algebra and Geometry, 3 (2):231–255, ???? 2019. CO-DEN SJAABQ. ISSN 2470-6566.

Chen:2017:FSS

[CGY17] Jingmin Chen, Sara Grundel, and Thomas P. Y. Yu. A flexible C^{2} subdivision scheme on the sphere: With application to biomembrane modelling. SIAM Journal on Applied Algebra and Geometry, 1(1):459–483, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Christandl:2023:GST

[CGZ23] Matthias Christandl, Fulvio Gesmundo, and Jeroen Zuiddam. A gap in the subrank of tensors. SIAM Journal on Applied Algebra and Geometry, 7(4):742–767, November 2023. ISSN 2470-6566.

Coss:2018:LCE

[CHHM18] Owen Coss, Jonathan D. Hauenstein, Hoon Hong, and Daniel K. Molzahn. Locating and counting equilibria of the Kuramoto model with rank-one coupling. SIAM Journal on Applied Algebra and Geometry, 2(1):45–71, ???? 2018. CODEN SJAABQ. ISSN 2470-6566.

Chindris:2022:QIT

[CI22] Calin Chindris and Jasim Ismaeel. A quiver invariant theoretic approach to radial isotropy and the paulsen problem for matrix frames. SIAM Journal on Applied Algebra and Geometry, 6(4):536– 562, ???? 2022. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/21M141470X.

Chistikov:2017:NMF

[CKM^{+}17] Dmitry Chistikov, Stefan Kiefer, Ines Marušić, Mahsa Shirmohammadi, and James Worrell. Nonnegative matrix factorization requires irrationality. SIAM Journal on Applied Algebra and Geometry, 1(1):285–307, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Cai:2021:ERS

[CKMW21] Chen Cai, Woojin Kim, Facundo Mémoli, and Yusu Wang. Elder-rule-staircodes for augmented metric spaces. SIAM Journal on Applied Algebra and Geometry, 5(3):417– 454, ???? 2021. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/20M1353605.

CampsMoreno:2024:ENS

[CLM24] Eduardo Camps Moreno, Hiram H. López, and Gretchen L. Matthews. Explicit non-special divisors of small degree, algebraic geometric hulls, and LCD codes from Kummer extensions.

SIAM Journal on Applied Algebra and Geometry, 8(2):394– 413, ???? 2024. CODEN SJAABQ. ISSN 2470-6566.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/21M1467936.

Cazzaniga:2024:WPR

[CLR24] Alberto Cazzaniga, Antonio Lerario, and Andrea Rosana. What is the probability that a random symmetric tensor is close to rank-one? SIAM Journal on Applied Algebra and Geometry, 8(2):227–258, April 2024. ISSN 2470-6566.

Chen:2017:ASR

[CLS17] Hao Chen, Kristin Lauter, and Katherine E. Stange. Attacks on the search RLWE problem with small errors. SIAM Journal on Applied Algebra and Geometry, 1(1):665–682, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Coste:2020:RLS

[CM20] Michel Coste and Seydou Moussa. Rationality of the locus of singularities of the general Gough–Stewart platform. SIAM Journal on Applied Algebra and Geometry, 4(3):401–421, ???? 2020. CODEN SJAABQ. ISSN 2470-6566.

Coons:2023:SCG

[CMMS23] Jane I. Coons, Aida Maraj, Pratik Misra, and Miruna-Stefana Sorea. Symmetrically colored Gaussian graphical models with toric vanishing ideals.

SIAM Journal on Applied Algebra and Geometry, 7(1):133– 158, ???? 2023. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/21M1466943.

Cahill:2017:CIA

[CMS17] Jameson Cahill, Dustin G.

Mixon, and Nate Strawn. Connectivity and irreducibility of algebraic varieties of finite unit norm tight frames. SIAM Journal on Applied Algebra and Geometry, 1(1):38–72, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Carlsson:2021:RHC

[CMS21] Gunnar Carlsson, Facundo Mémoli, and Santiago Segarra. Robust hierarchical clustering for directed networks: An axiomatic approach. SIAM Journal on Applied Algebra and Geometry, 5(4):675–700, ???? 2021. CODEN SJAABQ. ISSN 2470-6566. URL https://

epubs.siam.org/doi/10.1137/
20M1359201.

Carbone:2020:ESR

[CNN20] Lisa Carbone, Vidit Nanda, and Yusra Naqvi. Equivariant simplicial reconstruction. SIAM Journal on Applied Algebra and Geometry, 4(4):532–552, ???? 2020. CODEN SJAABQ. ISSN 2470-6566.

Cifuentes:2017:CNP

[CP17] Diego Cifuentes and Pablo A.

Parrilo. Chordal networks of polynomial ideals. SIAM Journal on Applied Algebra and Geometry, 1(1):73–110, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Comon:2017:IRD

[CQU17] Pierre Comon, Yang Qi, and Konstantin Usevich. Identifiability of an X-rank decomposition of polynomial maps. SIAM Journal on Applied Algebra and Geometry, 1(1):388–414, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Chacholski:2020:MSO

[CR20] Wojciech Chachólski and Henri Riihimäki. Metrics and stabilization in one parameter persistence. SIAM Journal on Applied Algebra and Geometry, 4 (1):69–98, ???? 2020. CODEN SJAABQ. ISSN 2470-6566.

Craciun:2019:PDS

[Cra19] Gheorghe Craciun. Polynomial dynamical systems, reaction networks, and toric differential inclusions. SIAM Journal on Applied Algebra and Geometry, 3 (1):87–106, ???? 2019. CODEN SJAABQ. ISSN 2470-6566.

Cantarella:2022:CCB

[CS22] Jason Cantarella and Henrik Schumacher. Computing the conformal barycenter. SIAM Journal on Applied Algebra and Geometry, 6(3):503– 530, ???? 2022. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/21M1449282.

DiRocco:2023:CGF

[DEE^{+}23] Sandra Di Rocco, Parker B. Edwards, David Eklund, Oliver Gäfvert, and Jonathan D. Hauenstein. Computing geometric feature sizes for algebraic manifolds. SIAM Journal on Applied Algebra and Geometry, 7(4):716–741, November 2023. ISSN 2470-6566.

Diehl:2022:TTS

[DEFT22] Joscha Diehl, Kurusch Ebrahimi-Fard, and Nikolas Tapia. Tropical time series, iterated-sums signatures, and quasisymmetric functions. SIAM Journal on Applied Algebra and Geometry, 6 (4):563–599, ???? 2022. CO-DEN SJAABQ. ISSN 2470-

6566. URL https://epubs.
siam.org/doi/10.1137/20M1380041.

Derksen:2018:GTS

[Der18] Harm Derksen. A general theory of singular values with applications to signal denoising. SIAM Journal on Applied Algebra and Geometry, 2(4):535–596, ???? 2018. CODEN SJAABQ. ISSN 2470-6566.

DiRocco:2020:BDA

[DEW20] Sandra Di Rocco, David Eklund, and Madeleine Weinstein. The bottleneck degree of algebraic varieties. SIAM Journal on Applied Algebra and Geometry, 4(1):227–253, ???? 2020. CODEN SJAABQ. ISSN 2470-

Dimitrova:2024:AED

[DFR^{+}24] Elena S. Dimitrova, Cameron H. Fredrickson, Nicholas A. Rondoni, Brandilyn Stigler, and Alan Veliz-Cuba. Algebraic experimental design: Theory and computation. SIAM Journal on Applied Algebra and Geometry, 8(2):284–301, April 2024. ISSN 2470-6566.

Dong:2023:DED

[DGHP23] Ruiwen Dong, Christian Goodbrake, Heather A. Harrington, and Gleb Pogudin. Differential elimination for dynamical models via projections with applications to structural identifiability. SIAM Journal on Applied Algebra and Geometry, 7(1): 194–235, ???? 2023. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/22M1469067.

DiRocco:2024:GLG

[DGS24] Sandra Di Rocco, Lukas Gustafsson, and Luca Schaffler. Gaussian likelihood geometry of projective varieties. SIAM Journal on Applied Algebra and Geometry, 8(1):89–113, February 2024. ISSN 2470-6566.

Deng:2023:FSC

[DH23] Yiyang Deng and Marshall Hampton. Finiteness of spatial central configurations with fixed subconfigurations. SIAM Journal on Applied Algebra and Geometry, 7(3):610– 622, ???? 2023. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/22M1490788.

Dressler:2017:PSN

[DIdW17] Mareike Dressler, Sadik Iliman, and Timo de Wolff. A Positivstellensatz for sums of nonnegative circuit polynomials. SIAM Journal on Applied Algebra and Geometry, 1(1):536–555, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Dorfler:2020:GCT

[DIP20] Julian Dörfler, Christian Ikenmeyer, and Greta Panova. On geometric complexity theory: Multiplicity obstructions are stronger than occurrence obstructions. SIAM Journal on Applied Algebra and Geometry, 4(2):354–376, ???? 2020. CO-DEN SJAABQ. ISSN 2470-6566.

Dostert:2023:LPF

[DJ23] Maria Dostert and Katharina Jochemko. Learning polytopes with fixed facet directions. SIAM Journal on Applied Algebra and Geometry, 7(2):440– 469, ???? 2023. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/22M1481695.

Duff:2022:GMG

[DKPR22] Timothy Duff, Viktor Korotynskiy, Tomas Pajdla, and Margaret H. Regan. Galois/Monodromy groups for decomposing minimal problems in 3D reconstruction.

SIAM Journal on Applied Algebra and Geometry, 6(4):740– 772, ???? 2022. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/21M1422872.

DeFord:2019:TVI

[DLSS19] Daryl DeFord, Hugo Lavenant, Zachary Schutzman, and Justin Solomon. Total variation isoperimetric profiles. SIAM Journal on Applied Algebra and Geometry, 3(4):585–613, ???? 2019. CODEN SJAABQ. ISSN 2470-6566.

Duursma:2021:MAD

[DLW21] Iwan Duursma, Xiao Li, and Hsin-Po Wang. Multilinear algebra for distributed storage. SIAM Journal on Applied Algebra and Geometry, 5(3):552– 587, ???? 2021. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/20M1346742.

Derksen:2021:MLE

[DM21] Harm Derksen and Visu Makam. Maximum likelihood estimation for matrix normal models via quiver representations. SIAM Journal on Applied Algebra and Geometry, 5(2):338– 365, ???? 2021. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/20M1369348.

Dressler:2022:APS

[DM22] Mareike Dressler and Riley Murray. Algebraic perspectives on signomial optimization.

SIAM Journal on Applied Algebra and Geometry, 6(4):650– 684, ???? 2022. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/21M1462568.

Devadoss:2017:SPN

[DP17] Satyan L. Devadoss and Samantha Petti. A space of phylogenetic networks. SIAM Journal on Applied Algebra and Geometry, 1(1):683–705, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Draisma:2024:ECC

[Dra24] Jan Draisma. Erratum: A counterexample to Comon’s conjecture. SIAM Journal on Applied Algebra and Geometry, 8(1):225, March 2024. ISSN 2470-6566. See [Shi18].

Dixon:2020:MMG

[DS20] Kael Dixon and Simon Salamon.

Moment maps and Galois orbits in quantum information theory. SIAM Journal on Applied Algebra and Geometry, 4(4):502–531, ???? 2020. CODEN SJAABQ. ISSN 2470-6566.

DiPasquale:2021:LBS

[DV21] Michael DiPasquale and Nelly Villamizar. A lower bound for splines on tetrahedral vertex stars. SIAM Journal on Applied Algebra and Geometry, 5(2): 250–277, ???? 2021. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/20M1341118.

DiPasquale:2024:BST

[DY24] Michael DiPasquale and Beihui Yuan. Bivariate splines on a triangulation with a single totally interior edge. SIAM Journal on Applied Algebra and Geometry, 8(3):686– 712, ???? 2024. CODEN SJAABQ. ISSN 2470-6566.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/23M158317X.

Dym:2018:ERS

[Dym18] Nadav Dym. Exact recovery with symmetries for the doubly stochastic relaxation. SIAM Journal on Applied Algebra and Geometry, 2(3):462–488, ???? 2018. CODEN SJAABQ. ISSN 2470-6566.

Edidin:2019:GAO

[Edi19] Dan Edidin. The geometry of ambiguity in one-dimensional phase retrieval. SIAM Journal on Applied Algebra and Geometry, 3(4):644–660, ???? 2019. CODEN SJAABQ. ISSN 2470-6566.

ElKhadir:2020:SSR

[El 20] Bachir El Khadir. On sum of squares representation of convex forms and generalized Cauchy– Schwarz inequalities. SIAM Journal on Applied Algebra and Geometry, 4(2):377–400, ???? 2020. CODEN SJAABQ. ISSN 2470-6566.

Forsgaard:2022:ABS

[FdW22] Jens Forsgård and Timo de Wolff. The algebraic boundary of the sonc-cone. SIAM Journal on Applied Algebra and Geometry, 6(3):468–502, ???? 2022. CO-DEN SJAABQ. ISSN 2470-

6566. URL https://epubs.
siam.org/doi/10.1137/20M1325484.

Fawzi:2018:LBP

[FE18] Hamza Fawzi and Mohab Safey El Din. A lower bound on the positive semidefinite rank of convex bodies. SIAM Journal on Applied Algebra and Geometry, 2(1):126–139, ???? 2018. CO-DEN SJAABQ. ISSN 2470-6566.

Freij-Hollanti:2017:PIR

[FHGHK17] Ragnar Freij-Hollanti, Oliver W.

Gnilke, Camilla Hollanti, and David A. Karpuk. Private information retrieval from coded databases with colluding servers. SIAM Journal on Applied Algebra and Geometry, 1(1):647–664, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Fickus:2018:PRP

[FJM18] Matthew Fickus, John Jasper, and Dustin G. Mixon. Packings in real projective spaces. SIAM Journal on Applied Algebra and Geometry, 2(3):377–409, ???? 2018. CODEN SJAABQ. ISSN 2470-6566.

Fiore:2018:VTT

[FN18] Thomas M. Fiore and Thomas Noll. Voicing transformations of triads. SIAM Journal on Applied Algebra and Geometry, 2 (2):281–313, ???? 2018. CO-DEN SJAABQ. ISSN 2470-6566.

Fritz:2017:SCO

[FNT17] Tobias Fritz, Tim Netzer, and Andreas Thom. Spectrahedral containment and operator systems with finite-dimensional realization. SIAM Journal on Applied Algebra and Geometry, 1 (1):556–574, ???? 2017. CO-DEN SJAABQ. ISSN 2470-6566.

Faulstich:2024:CCT

[FO24] Fabian M. Faulstich and Mathias Oster. Coupled cluster theory: Toward an algebraic geometry formulation. SIAM Journal on Applied Algebra and Geometry, 8(1):138–188, February 2024. ISSN 2470-6566.

Forcey:2023:CPE

[For23] Stefan Forcey. Circular planar electrical networks, split systems, and phylogenetic networks. SIAM Journal on Applied Algebra and Geometry, 7 (1):49–76, ???? 2023. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/22M1473844.

Fritz:2023:AVP

[Fri23] Tobias Fritz. Abstract Vergleichsstellensätze for preordered semifields and semirings I. SIAM Journal on Applied Algebra and Geometry, 7(2):505– 547, ???? 2023. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/22M1498413.

Goodman:2024:RSO

[GC24] Jacob R. Goodman and Leonardo J.

Colombo. Reduction by symmetry in obstacle avoidance problems on Riemannian manifolds. SIAM Journal on Applied Algebra and Geometry, 8(1):26–53, January 2024. ISSN 2470-6566.

Goucha:2021:PRM

[GG21] António Pedro Goucha and João Gouveia. The phaseless rank of a matrix. SIAM Journal on Applied Algebra and Geometry, 5 (3):526–551, ???? 2021. CO-DEN SJAABQ. ISSN 2470-

6566. URL https://epubs.
siam.org/doi/10.1137/19M1289820.

Ghrist:2017:PAD

[GK17] Robert Ghrist and Sanjeevi Krishnan. Positive Alexander duality for pursuit and evasion. SIAM Journal on Applied Algebra and Geometry, 1(1):308–327, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Gross:2018:DPN

[GL18] Elizabeth Gross and Colby Long. Distinguishing phylogenetic networks. SIAM Journal on Applied Algebra and Geometry, 2(1):72–93, ???? 2018. CODEN SJAABQ. ISSN 2470-6566.

Grigsby:2022:TBH

[GL22] J. Elisenda Grigsby and Kathryn Lindsey. On transversality of bent hyperplane arrangements and the topological expressiveness of ReLU neural networks.

SIAM Journal on Applied Algebra and Geometry, 6(2):216– 242, ???? 2022. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/20M1368902.

Giusti:2023:SLF

[GL23] Chad Giusti and Darrick Lee. Signatures, Lipschitz-free spaces, and paths of persistence diagrams. SIAM Journal on Applied Algebra and Geometry, 7 (4):828–866, December 2023. ISSN 2470-6566.

Gallet:2019:RSO

[GLSV19] Matteo Gallet, Niels Lubbes, Josef Schicho, and Jan Vrsek. Reconstruction of surfaces with ordinary singularities from their silhouettes. SIAM Journal on Applied Algebra and Geometry, 3(3):472–506, ???? 2019. CO-DEN SJAABQ. ISSN 2470-6566.

Grinberg:2019:BMD

[GO19a] Darij Grinberg and Peter J. Olver. The n body matrix and its determinant. SIAM Journal on Applied Algebra and Geometry, 3(1):67–86, ???? 2019. CODEN SJAABQ. ISSN 2470-6566.

Grindstaff:2019:RPL

[GO19b] Gillian Grindstaff and Megan Owen. Representations of partial leaf sets in phylogenetic tree space. SIAM Journal on Applied Algebra and Geometry, 3 (4):691–720, ???? 2019. CO-DEN SJAABQ. ISSN 2470-6566.

Gruica:2022:CCL

[GR22] Anina Gruica and Alberto Ravagnani. Common complements of linear subspaces and the sparseness of MRD codes. SIAM Journal on Applied Algebra and Geometry, 6(2):79– 110, ???? 2022. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/21M1428947.

Himmelmann:2024:RGM

[HE24] Matthias Himmelmann and Myfanwy E. Evans. Robust geometric modeling of 3-periodic tensegrity frameworks using Riemannian optimization. SIAM Journal on Applied Algebra and Geometry, 8(2):259–283, April 2024. ISSN 2470-6566.

Hamada:2021:CNR

[HH21] Masaki Hamada and Hiroshi Hirai. Computing the ncrank via discrete convex optimization on CAT(0) spaces. SIAM Journal on Applied Algebra and Geometry, 5(3):455– 478, ???? 2021. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/20M138836X.

Heaton:2021:SAG

[HHS21] Alexander Heaton, Serkan Hosten, and Isabelle Shankar. Symmetry adapted Gram spectrahedra. SIAM Journal on Applied Algebra and Geometry, 5 (1):140–164, ???? 2021. CO-DEN SJAABQ. ISSN 2470-

                                        6566. URL https://epubs.
                                        siam.org/doi/10.1137/20M133796X.

Hirai:2019:CDD

[Hir19] Hiroshi Hirai. Computing the degree of determinants via discrete convex optimization on Euclidean buildings. SIAM Journal on Applied Algebra and Geometry, 3(3):523–557, ???? 2019. CODEN SJAABQ. ISSN 2470-6566.

Hackbusch:2017:PHO

[HKU17] Wolfgang Hackbusch, Daniel Kressner, and André Uschmajew. Perturbation of higherorder singular values. SIAM Journal on Applied Algebra and Geometry, 1(1):374–387, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Hayut:2023:CSI

[HL23] Yair Hayut and David Lehavi. Complete SE(3) invariants for a comeagre set of C^{3} compact orientable surfaces in R^{3}. SIAM Journal on Applied Algebra and Geometry, 7(2):311– 344, ???? 2023. CODEN SJAABQ. ISSN 2470-6566.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/21M1445776.

Han:2022:NBW

[HM22] Kangjin Han and Hyunsuk Moon. A new bound for the Waring rank of monomials. SIAM Journal on Applied Algebra and Geometry, 6(3):

407–431, ???? 2022. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/21M1390736.

Hubert:2024:OSW

[HMR24] Evelyne Hubert, Tobias Metzlaff, and Cordian Riener. Orbit spaces of Weyl groups acting on compact tori: A unified and explicit polynomial description. SIAM Journal on Applied Algebra and Geometry, 8(3): 612–649, ???? 2024. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/23M158173X.

Helmer:2023:CLH

[HN23] Martin Helmer and Vidit Nanda. Complex links and Hilbert-samuel multiplicities. SIAM Journal on Applied Algebra and Geometry, 7(1):29– 48, ???? 2023. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/22M1475533.

Harrington:2019:SMP

[HOST19] Heather A. Harrington, Nina Otter, Hal Schenck, and Ulrike Tillmann. Stratifying multiparameter persistent homology. SIAM Journal on Applied Algebra and Geometry, 3(3):439–471, ???? 2019. CODEN SJAABQ. ISSN 2470-6566.

Hauenstein:2018:ESG

[HSW18] Jonathan D. Hauenstein, Samantha N. Sherman, and Charles W. Wampler. Exceptional Stewart– Gough platforms, Segrè embeddings, and the special Euclidean group. SIAM Journal on Applied Algebra and Geometry, 2 (1):179–205, ???? 2018. CO-DEN SJAABQ. ISSN 2470-6566.

Ito:2023:ARO

[IL23] Masaru Ito and Bruno F. Lourenço. Automorphisms of rank-one generated hyperbolicity cones and their derivative relaxations. SIAM Journal on Applied Algebra and Geometry, 7 (1):236–263, ???? 2023. CO-DEN SJAABQ. ISSN 2470-

6566. URL https://epubs.
siam.org/doi/10.1137/22M1513964.

Ilten:2017:FST

[IZ17] Nathan Ilten and Alexandre Zotine. On Fano schemes of toric varieties. SIAM Journal on Applied Algebra and Geometry, 1 (1):152–174, ???? 2017. CO-DEN SJAABQ. ISSN 2470-6566.

Jany:2023:PMM

[Jan23]  Benjamin Jany.  The pro-

jectivization matroid of a qmatroid. SIAM Journal on Applied Algebra and Geometry, 7 (2):386–413, ???? 2023. CO-DEN SJAABQ. ISSN 2470-

6566. URL https://epubs.
siam.org/doi/10.1137/22M1494567.

Jeffs:2020:MNC

[Jef20] R. Amzi Jeffs. Morphisms of neural codes. SIAM Journal on Applied Algebra and Geometry, 4(1):99–122, ???? 2020.

CODEN SJAABQ. ISSN 2470-6566. See erratum [Jef22].

Jeffs:2022:EMN

[Jef22] R. Amzi Jeffs. Erratum for “Morphisms of Neural Codes”. SIAM Journal on Applied Algebra and Geometry, 6(1):77– 78, ???? 2022. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/21M1450690.

See [Jef20].

Joswig:2022:GPO

[JKS22] Michael Joswig, Max Klimm, and Sylvain Spitz. Generalized permutahedra and optimal auctions. SIAM Journal on Applied Algebra and Geometry, 6(4): 711–739, ???? 2022. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/21M1441286.

Javaloyes:2023:GMW

[JPRS23] Miguel Ángel Javaloyes, Enrique Pendás-Recondo, and Miguel Sánchez. A general model for wildfire propagation with wind and slope. SIAM Journal on Applied Algebra and Geometry, 7(2):414– 439, ???? 2023. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/22M1477866.

Kolbe:2022:EIC

[KE22] Benedikt Kolbe and Myfanwy E.

Evans. Enumerating isotopy classes of tilings guided by the symmetry of triply periodic minimal surfaces. SIAM Journal on Applied Algebra and Geometry, 6(1):1–40, ???? 2022. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/20M1358943.

Kileel:2017:MPC

[Kil17] Joe Kileel. Minimal problems for the calibrated trifocal variety. SIAM Journal on Applied Algebra and Geometry, 1(1):575–598, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Kech:2017:OIC

[KK17] Michael Kech and Felix Krahmer. Optimal injectivity conditions for bilinear inverse problems with applications to identifiability of deconvolution problems. SIAM Journal on Applied Algebra and Geometry, 1 (1):20–37, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Kahle:2017:GRO

[KKKR17] Thomas Kahle, Kaie Kubjas, Mario Kummer, and Zvi Rosen. The geometry of rank-one tensor completion. SIAM Journal on Applied Algebra and Geometry, 1(1):200–221, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Kulkarni:2021:AIG

[KL21] Avinash Kulkarni and Antonio Lerario. p-adic integral geometry. SIAM Journal on Applied Algebra and Geometry, 5 (1):28–59, ???? 2021. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/19M1284737.

Kozhasov:2023:VSS

[KL23] Khazhgali Kozhasov and Jean B.

Lasserre. Volumes of sublevel sets of nonnegative forms and complete monotonicity. SIAM Journal on Applied Algebra and Geometry, 7(4):768–785, November 2023. ISSN 2470-6566.

Kaveh:2019:KBH

[KM19] Kiumars Kaveh and Christopher Manon. Khovanskii bases, higher rank valuations, and tropical geometry. SIAM Journal on Applied Algebra and Geometry, 3(2):292–336, ???? 2019. CODEN SJAABQ. ISSN 2470-6566.

Kohn:2022:GLC

[KMMT22] Kathlén Kohn, Thomas Merkh, Guido Montúfar, and Matthew Trager. Geometry of linear convolutional networks. SIAM Journal on Applied Algebra and Geometry, 6(3):368– 406, ???? 2022. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/21M1441183.

Kohn:2024:FSC

[KMST24] Kathlén Kohn, Guido Montúfar, Vahid Shahverdi, and Matthew Trager. Function space and critical points of linear convolutional networks. SIAM Journal on Applied Algebra and Geometry, 8 (2):333–362, May 2024. ISSN 2470-6566.

Kepley:2021:CLE

[KMZ21] Shane Kepley, Konstantin Mischaikow, and Lun Zhang. Computing linear extensions for polynomial posets subject to algebraic constraints. SIAM Journal on Applied Algebra and Geometry, 5(2):388– 416, ???? 2021. CODEN SJAABQ. ISSN 2470-6566.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/20M1343208.

Kopp:2017:AGR

[Kop17] Gene S. Kopp. The arithmetic geometry of resonant Rossby wave triads. SIAM Journal on Applied Algebra and Geometry, 1(1):352–373, ???? 2017. CO-DEN SJAABQ. ISSN 2470-6566.

Koshelev:2022:OEE

[Kos22] Dmitrii Koshelev. Optimal encodings to elliptic curves of j-invariants 0, 1728. SIAM Journal on Applied Algebra and Geometry, 6(4):600–617, ???? 2022. CODEN SJAABQ. ISSN 2470-6566. URL https://

                                        epubs.siam.org/doi/10.1137/
                                        21M1441602.

Kao:2023:FTL

[KOT23] Chiu-Yen Kao, Braxton Osting, and Jackson C. Turner. Flat tori with large Laplacian eigenvalues in dimensions up to eight. SIAM Journal on Applied Algebra and Geometry, 7(1):172– 193, ???? 2023. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/22M1478823.

Kozhasov:2018:FRE

[Koz18] Khazhgali Kozhasov. On fully real eigenconfigurations of tensors. SIAM Journal on Applied Algebra and Geometry, 2 (2):339–347, ???? 2018. CO-DEN SJAABQ. ISSN 2470-6566.

Kravaris:2023:DEP

[Kra23] Cosmas Kravaris. On the density of eigenvalues on periodic graphs. SIAM Journal on Applied Algebra and Geometry, 7 (3):585–609, ???? 2023. CO-DEN SJAABQ. ISSN 2470-

6566. URL https://epubs.
siam.org/doi/10.1137/22M1505268.

Kogan:2020:DSA

[KRV20] Irina A. Kogan, Michael Ruddy, and Cynthia Vinzant. Differential signatures of algebraic curves. SIAM Journal on Applied Algebra and Geometry, 4 (1):185–226, ???? 2020. CO-DEN SJAABQ. ISSN 2470-6566.

Klep:2024:RCM

[KSZ24] Igor Klep, Tea Strekelj, and Aljaz Zalar. A random copositive matrix is completely positive with positive probability. SIAM Journal on Applied Algebra and Geometry, 8(3):583– 611, ???? 2024. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/23M1583491.

Koutsoumpelias:2020:MIL

[KW20] Alexandros Grosdos Koutsoumpelias and Markus Wageringel. Moment ideals of local Dirac mixtures. SIAM Journal on Applied Algebra and Geometry, 4(1):1–27, ???? 2020. CODEN SJAABQ. ISSN 2470-6566.

Knudsen:2024:ACC

[KW24] Ben Knudsen and Shmuel Weinberger. Analog category and complexity. SIAM Journal on Applied Algebra and Geometry, 8(3):713–732, ???? 2024. CO-DEN SJAABQ. ISSN 2470-

                                        6566. URL https://epubs.
                                        siam.org/doi/10.1137/24M1635004.

Lasserre:2019:VSS

[Las19] Jean B. Lasserre. Volume of sublevel sets of homogeneous polynomials. SIAM Journal on Applied Algebra and Geometry, 3 (2):372–389, ???? 2019. CO-DEN SJAABQ. ISSN 2470-6566.

Landsberg:2017:GBR

[LM17] J. M. Landsberg and Mateusz Michałek. On the geometry of border rank decompositions for matrix multiplication and other tensors with symmetry. SIAM Journal on Applied Algebra and Geometry, 1(1):2–19, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Lasserre:2020:CHB

[LM20] Jean B. Lasserre and Victor Magron. Computing the Hausdorff boundary measure of semialgebraic sets. SIAM Journal on Applied Algebra and Geometry, 4 (3):441–469, ???? 2020. CO-DEN SJAABQ. ISSN 2470-6566.

Laface:2023:DAT

[LMR23] Antonio Laface, Alex Massarenti, and Rick Rischter. Decomposition algorithms for tensors and polynomials. SIAM Journal on Applied Algebra and Geometry, 7(1):264–290, ???? 2023. CODEN SJAABQ. ISSN 2470-

6566. URL https://epubs.
siam.org/doi/10.1137/21M1453712.

Lindberg:2023:MLD

[LNRW23] Julia Lindberg, Nathan Nicholson, Jose I. Rodriguez, and Zinan Wang. The maximum likelihood degree of sparse polynomial systems. SIAM Journal on Applied Algebra and Geometry, 7(1):159– 171, ???? 2023. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/21M1422550.

Lim:2024:TSD

[LON24] Uzu Lim, Harald Oberhauser, and Vidit Nanda. Tangent space and dimension estimation with the Wasserstein distance. SIAM Journal on Applied Algebra and Geometry, 8(3):650– 685, ???? 2024. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/22M1522711.

Le:2023:TCM

[LR23] Dinh V. Le and Tim Römer. Theorems of Carathéodory, Minkowski–Weyl, and Gordan up to symmetry. SIAM Journal on Applied Algebra and Geometry, 7(1):291–310, ???? 2023. CODEN SJAABQ. ISSN 2470-

                                        6566. URL https://epubs.
                                        siam.org/doi/10.1137/22M148865X.

Lesnick:2020:QGI

[LRR20] Michael Lesnick, Raúl Rabadán, and Daniel I. S. Rosenbloom. Quantifying genetic innovation: Mathematical foundations for the topological study of reticulate evolution. SIAM Journal on Applied Algebra and Geometry, 4(1):141–184, ???? 2020. CODEN SJAABQ. ISSN 2470-6566.

Ladisch:2018:ELO

[LS18a] Frieder Ladisch and Achill Schürmann. Equivalence of lattice orbit polytopes. SIAM Journal on Applied Algebra and Geometry, 2(2):259–280, ???? 2018. CODEN SJAABQ. ISSN 2470-6566.

Little:2018:CSS

[LS18b] John Little and Hal Schenck. Codes from surfaces with small Picard number. SIAM Journal on Applied Algebra and Geometry, 2(2):242–258, ???? 2018. CODEN SJAABQ. ISSN 2470-6566.

Lairez:2019:NTM

[LS19] Pierre Lairez and Emre Can Sertöz. A numerical transcendental method in algebraic geometry: Computation of Picard groups and related invariants. SIAM Journal on Applied Algebra and Geometry, 3(4):559–584, ???? 2019. CODEN SJAABQ. ISSN 2470-6566.

Lieblich:2020:THS

[LV20] Max Lieblich and Lucas Van Meter. Two Hilbert schemes in computer vision. SIAM Journal on Applied Algebra and Geometry, 4(2):297–321, ???? 2020. CODEN SJAABQ. ISSN 2470-6566.

Lesnick:2022:CMP

[LW22] Michael Lesnick and Matthew Wright. Computing minimal presentations and bigraded Betti numbers of 2parameter persistent homology. SIAM Journal on Applied Algebra and Geometry, 6(2):267– 298, ???? 2022. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/20M1388425.

Miolane:2018:TCT

[MHP18] Nina Miolane, Susan Holmes, and Xavier Pennec. Topologically constrained template estimation via Morse–Smale complexes controls its statistical consistency. SIAM Journal on Applied Algebra and Geometry, 2(2):348–375, ???? 2018. CO-DEN SJAABQ. ISSN 2470-6566.

Muller:2019:BFE

[MHR19] Stefan Müller, Josef Hofbauer, and Georg Regensburger. On the bijectivity of families of exponential/generalized polynomial maps. SIAM Journal on Applied Algebra and Geometry, 3(3):412–438, ???? 2019. CO-DEN SJAABQ. ISSN 2470-6566.

Micheli:2019:SPD

[Mic19] Giacomo Micheli. On the selection of polynomials for the DLP quasi-polynomial time algorithm for finite fields of small characteristic. SIAM Journal on Applied Algebra and Geometry, 3(2):256–265, ???? 2019. CODEN SJAABQ. ISSN 2470-6566.

Monod:2019:TSS

[MKPGC19] Anthea Monod, Sara Kalisnik, Juan Ángel Patiño-Galindo, and Lorin Crawford. Tropical sufficient statistics for persistent homology. SIAM Journal on Applied Algebra and Geometry, 3 (2):337–371, ???? 2019. CO-DEN SJAABQ. ISSN 2470-6566.

Munthe-Kaas:2020:ICL

[MKSV20] Hans Z. Munthe-Kaas, Ari Stern, and Olivier Verdier. Invariant connections, Lie algebra actions, and foundations of numerical integration on manifolds. SIAM Journal on Applied Algebra and Geometry, 4 (1):49–68, ???? 2020. CODEN SJAABQ. ISSN 2470-6566.

Montufar:2017:DMK

[MM17] Guido Montúfar and Jason Morton. Dimension of marginals of Kronecker product models. SIAM Journal on Applied Algebra and Geometry, 1(1):126–151, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Michalek:2021:MLD

[MMW21] Mateusz Michałek, Leonid Monin, and Jarosław A. Wiśniewski. Maximum likelihood degree, complete quadrics, and C^{∗}- action. SIAM Journal on Applied Algebra and Geometry, 5 (1):60–85, ???? 2021. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/20M1335960.

McCleary:2022:EDP

[MP22] Alexander McCleary and Amit Patel. Edit distance and persistence diagrams over lattices. SIAM Journal on Applied Algebra and Geometry, 6(2):134– 155, ???? 2022. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/20M1373700.

Maxim:2020:EDD

[MRW20] Laurentiu G. Maxim, Jose I. Rodriguez, and Botong Wang. Euclidean distance degree of the multiview variety. SIAM Journal on Applied Algebra and Geometry, 4(1):28–48, ???? 2020. CODEN SJAABQ. ISSN 2470-

Montufar:2022:SBN

[MRZ22] Guido Montúfar, Yue Ren, and Leon Zhang. Sharp bounds for the number of regions of maxout networks and vertices of Minkowski sums. SIAM Journal on Applied Algebra and Geometry, 6(4):618– 649, ???? 2022. CODEN SJAABQ. ISSN 2470-6566.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/21M1413699.

Meunier:2019:MVS

[MS19] Frédéric Meunier and Francis Edward Su. Multilabeled versions of Sperner’s and Fan’s lemmas and applications. SIAM Journal on Applied Algebra and Geometry, 3(3):391–411, ???? 2019. CODEN SJAABQ. ISSN 2470-6566.

Malic:2023:CCP

[MS23] Goran Malić and Ileana Streinu. Computing circuit polynomials in the algebraic rigidity matroid. SIAM Journal on Applied Algebra and Geometry, 7(2):345– 385, ???? 2023. CODEN SJAABQ. ISSN 2470-6566.

                                        URL https://epubs.siam.
                                        org/doi/10.1137/21M1437986.

Meyer:2022:ZST

[MSS22] Kyle Meyer, Ivan Soprunov, and Jenya Soprunova. F_{q}- zeros of sparse trivariate polynomials and toric 3-fold codes. SIAM Journal on Applied Algebra and Geometry, 6(3):432– 467, ???? 2022. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/21M1436890.

Mischaikow:2023:CCI

[MW23] Konstantin Mischaikow and Charles Weibel. Computing the Conley index: a cautionary tale. SIAM Journal on Applied Algebra and Geometry, 7(4):809– 827, December 2023. ISSN 2470-6566.

Neff:2019:EGM

[NFB19] Patrizio Neff, Andreas Fischle, and Lev Borisov. Explicit global minimization of the symmetrized Euclidean distance by a characterization of real matrices with symmetric square. SIAM Journal on Applied Algebra and Geometry, 3(1):31–43, ???? 2019. CODEN SJAABQ. ISSN 2470-6566.

Nicklasson:2023:TNT

[Nic23] Lisa Nicklasson. Toric and non-toric Bayesian networks. SIAM Journal on Applied Algebra and Geometry, 7(3):549– 566, ???? 2023. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/22M1515690.

Nie:2017:STN

[Nie17] Jiawang Nie. Symmetric tensor nuclear norms. SIAM Journal on Applied Algebra and Geometry, 1(1):599–625, ???? 2017. CODEN SJAABQ. ISSN 2470-

Napolitano:2024:CTA

[NPSZ24] Vito Napolitano, Olga Polverino, Paolo Santonastaso, and Ferdinando Zullo. Clubs and their applications. SIAM Journal on Applied Algebra and Geometry, 8(3):493–518, ???? 2024. CO-DEN SJAABQ. ISSN 2470-

6566. URL https://epubs.
siam.org/doi/10.1137/22M1523534.

Needham:2023:GAP

[NW23] Tom Needham and Thomas Weighill. Geometric averages of partitioned datasets. SIAM Journal on Applied Algebra and Geometry, 7(1):104– 132, ???? 2023. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/21M1436580.

Obayashi:2018:VOC

[Oba18] Ippei Obayashi. Volume-optimal cycle: Tightest representative cycle of a generator in persistent homology. SIAM Journal on Applied Algebra and Geometry, 2(4):508–534, ???? 2018. CODEN SJAABQ. ISSN 2470-6566.

Ovchinnikov:2022:MEP

[OPPS22] Alexey Ovchinnikov, Anand Pillay, Gleb Pogudin, and Thomas Scanlon. Multiexperiment parameter identifiability of ODEs and model theory. SIAM Journal on Applied Algebra and Geometry, 6(3): 339–367, ???? 2022. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/21M1389845.

Oertel:2020:DFR

[OPW20] Timm Oertel, Joseph Paat, and Robert Weismantel. The distributions of functions related to parametric integer optimization. SIAM Journal on Applied Algebra and Geometry, 4(3):422–440, ???? 2020. CODEN SJAABQ. ISSN 2470-6566.

Oudot:2024:SMB

[OS24] Steve Oudot and Luis Scoccola.

On the stability of multigraded Betti numbers and Hilbert functions. SIAM Journal on Applied Algebra and Geometry, 8(1):54– 88, January 2024. ISSN 2470-6566.

Ozyesil:2018:SCM

[ÖSS18] Onur Özyesil, Nir Sharon, and Amit Singer. Synchronization over Cartan motion groups via contraction. SIAM Journal on Applied Algebra and Geometry, 2(2):207–241, ???? 2018. CO-DEN SJAABQ. ISSN 2470-6566.

Polak:2022:SRO

[Pol22] Sven C. Polak. Symmetry reduction to optimize a graph-based polynomial from queueing theory. SIAM Journal on Applied Algebra and Geometry, 6(2): 243–266, ???? 2022. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/21M1413298.

Pryhuber:2022:ETV

[PST22] Andrew Pryhuber, Rainer Sinn, and Rekha R. Thomas. Existence of two view chiral reconstructions. SIAM Journal on Applied Algebra and Geometry, 6(1):41–76, ???? 2022. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/20M1381848.

Plaumann:2021:KTJ

[PSW21] Daniel Plaumann, Rainer Sinn, and Stephan Weis. Kippenhahn’s theorem for joint numerical ranges and quantum states. SIAM Journal on Applied Algebra and Geometry, 5 (1):86–113, ???? 2021. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/19M1286578.

Ravagnani:2022:WNC

[Rav22]  Alberto Ravagnani.  Whit-

ney numbers of combinatorial geometries and higher-weight Dowling lattices. SIAM Journal on Applied Algebra and Geometry, 6(2):156–189, ???? 2022. CODEN SJAABQ. ISSN 2470-

6566. URL https://epubs.
siam.org/doi/10.1137/20M1382635.

Robeva:2021:MTS

[RS21] Elina Robeva and Jean-Baptiste Seby. Multi-trek separation in linear structural equation models. SIAM Journal on Applied Algebra and Geometry, 5(2): 278–303, ???? 2021. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/20M1316470.

Ruiz:2024:MTP

[RS24] Luis Crespo Ruiz and Francisco Santos. Multitriangulations and tropical Pfaffians. SIAM Journal on Applied Algebra and Geometry, 8(2):302–332, April 2024. ISSN 2470-6566.

Rodriguez:2017:MLD

[RW17] Jose Israel Rodriguez and Botong Wang. The maximum likelihood degree of mixtures of independence models. SIAM Journal on Applied Algebra and Geometry, 1(1):484–506, ???? 2017. CODEN SJAABQ. ISSN 2470-6566.

Saunderson:2019:CPN

[Sau19] James Saunderson. Certifying polynomial nonnegativity via hyperbolic optimization. SIAM Journal on Applied Algebra and Geometry, 3(4):661–690, ???? 2019. CODEN SJAABQ. ISSN 2470-6566.

Scheiderer:2018:SRC

[Sch18a] Claus Scheiderer. Semidefinite representation for convex hulls of real algebraic curves. SIAM Journal on Applied Algebra and Geometry, 2(1):1–25, ???? 2018. CODEN SJAABQ. ISSN 2470-6566.

Scheiderer:2018:SS

[Sch18b] Claus Scheiderer. Spectrahedral shadows. SIAM Journal on Applied Algebra and Geometry, 2 (1):26–44, ???? 2018. CODEN SJAABQ. ISSN 2470-6566.

Scheiderer:2021:SOC

[Sch21]  Claus Scheiderer.  Second-

order cone representation for convex sets in the plane. SIAM Journal on Applied Algebra and Geometry, 5(1):114– 139, ???? 2021. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/20M133717X.

Shen:2017:SBR

[SG17] Li-Yong Shen and Ron Goldman. Strong μ-bases for rational tensor product surfaces and extraneous factors associated to bad base points and anomalies at infinity. SIAM Journal on Applied Algebra and Geometry, 1(1):328–351, ???? 2017. CO-DEN SJAABQ. ISSN 2470-6566.

Shankar:2019:CGD

[Sha19] Shiva Shankar. Corrigendum: The Generic Degree of Autonomy. SIAM Journal on Applied Algebra and Geometry, 3 (1):172–174, ???? 2019. CO-DEN SJAABQ. ISSN 2470-6566. See [SR18].

Shitov:2018:CCC

[Shi18] Yaroslav Shitov. A counterexample to Comon’s conjecture. SIAM Journal on Applied Algebra and Geometry, 2(3):428–443, ???? 2018. CODEN SJAABQ. ISSN 2470-6566. See erratum [Dra24].

Shitov:2021:WRX

[Shi21] Yaroslav Shitov. The Waring rank of the 3 x 3 permanent. SIAM Journal on Applied Algebra and Geometry, 5(4):701– 714, ???? 2021. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/20M1349254.

Shitov:2022:SCS

[Shi22] Yaroslav N. Shitov. The subrank of a complex symmetric tensor can exceed its symmetric subrank. SIAM Journal on Applied Algebra and Geometry, 6(4): 531–535, ???? 2022. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/21M1465494.

Shankar:2018:GDA

[SR18] Shiva Shankar and Paula Rocha.

The generic degree of autonomy. SIAM Journal on Applied Algebra and Geometry, 2(3):410–427, ???? 2018. CODEN SJAABQ. ISSN 2470-6566. See corrigendum [Sha19].

Stange:2021:AAS

[Sta21] Katherine E. Stange. Algebraic aspects of solving Ring-LWE, including ring-based improvements in the Blum– Kalai–Wasserman algorithm. SIAM Journal on Applied Algebra and Geometry, 5(2):366– 387, ???? 2021. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/19M1280442.

Sturmfels:2017:MEC

[Stu17] Bernd Sturmfels. Message from the Editor-in-Chief. SIAM Journal on Applied Algebra and Geometry, 1(1):1, ???? 2017. CO-DEN SJAABQ. ISSN 2470-6566.

Tokcan:2021:AMT

[TGND21] Neriman Tokcan, Jonathan Gryak, Kayvan Najarian, and Harm Derksen. Algebraic methods for tensor data. SIAM Journal on Applied Algebra and Geometry, 5(1):1–27, ???? 2021. CODEN SJAABQ. ISSN 2470-

                                        6566. URL https://epubs.
                                        siam.org/doi/10.1137/19M1272494.

Tran:2021:FMB

[Tra21] Ngoc Mai Tran. The finite matroid-based valuation conjecture is false. SIAM Journal on Applied Algebra and Geometry, 5(3):506–525, ???? 2021. CODEN SJAABQ. ISSN 2470-

                                        6566. URL https://epubs.
                                        siam.org/doi/10.1137/19M1304295.

Turatti:2022:TDT

[Tur22] Ettore Turatti. On tensors that are determined by their singular tuples. SIAM Journal on Applied Algebra and Geometry, 6 (2):319–338, ???? 2022. CO-DEN SJAABQ. ISSN 2470-

                                        6566. URL https://epubs.
                                        siam.org/doi/10.1137/21M1412980.

Volcic:2019:SNP

[Vol19] Jurij Volcic. Stable noncommutative polynomials and their determinantal representations. SIAM Journal on Applied Algebra and Geometry, 3(1):152–171, ???? 2019. CODEN SJAABQ. ISSN 2470-6566.

Wang:2022:NPC

[Wan22] Jie Wang. Nonnegative polynomials and circuit polynomials. SIAM Journal on Applied Algebra and Geometry, 6(2): 111–133, ???? 2022. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/20M1313969.

Ye:2017:CCE

[YL17] Ke Ye and Lek-Heng Lim. Cohomology of cryo-electron microscopy. SIAM Journal on Applied Algebra and Geometry, 1 (1):507–535, ???? 2017. CO-DEN SJAABQ. ISSN 2470-6566.

Yu:2021:OCA

[YMV21] Haizi Yu, Igor Mineyev, and Lav R. Varshney. Orbit computation for atomically generated subgroups of isometries of Z^{n}. SIAM Journal on Applied Algebra and Geometry, 5(3): 479–505, ???? 2021. CODEN SJAABQ. ISSN 2470-6566.

URL https://epubs.siam.
org/doi/10.1137/20M1375127.